OpenStudy (anonymous):

Find the exact value of the radical expression in simplest form. the square root of 3 times x + 3 the square root of 3 times x − 2 the square root of 3 times x − 10 −10 2 times x times the square root of 3 times x 2 times the square root of 3 times x − 10 10 − 2 times x times the square root of 3 times x

OpenStudy (anonymous):

\[\sqrt{3x}+3\sqrt{3x}-2\sqrt{3x}-10\] now can you solve????